Download
Monthly QA Report

Quality Assurance

The Oklahoma Mesonet’s quality-assurance (QA) consists of four principal components:
1) laboratory calibration,
2) on-site intercomparison,
3) automated QA, and
4) manual QA.

In the instrument laboratory, all sensors are calibrated to validate or improve upon factory calibrations.

On an annual basis, the accuracy of several types of sensors in the field is verified through comparison to calibrated reference sensors housed in a portable system.

To make a first pass through the 640,000+ observations that are received at the Mesonet each day, automated QA software includes numerous algorithms to evaluate the data received from remote stations.

Finally, the Mesonet’s QA meteorologists employ numerous manual techniques to complement automated QA including analysis of monthly statistics to detect sensor drift or bias (click on a month and year to the left to download a PDF version of a Monthly Report). In addition to detecting problematic sensors, the QA meteorologists trace the true start time of each problem so that appropriate data can be manually flagged as erroneous. The QA meteorologists are responsible for communicating problems to, and coordinating with, appropriate field technicians to ensure proper resolution (see a list of the current outstanding tickets below).

Current Trouble Tickets in the Oklahoma Mesonet
(For questions regarding these tickets, contact )

Station Variable Date Reported Problem Description
BYAR 229L Delta Temperature at 5 cm 2021-01-14 20:02:01 Starting temp = final temp, suggesting failed heater. Please replace.
WATO 229L Delta Temperature at 25 cm 2021-01-14 15:46:25 Both starting and final temperature reporting -7999. Both soil temperature and soil moisture report errant or missing values.
OKCE Rain since midnight UTC at 0.6m 2021-01-04 22:10:58 Primary gauge fails to record precip. Before problem began, gauge had history of very low drip tests. Please replace gauge and RG cable.
BEAV Relative Humidity at 1.5 m 2020-12-16 15:56:59 If HMP155 cable not replaced after wildfire, please replace sensor cabling.
BEAV Relative Humidity at 1.5 m 2020-12-16 15:55:05 RELH sometimes reports values >110%, then suddenly drops to 10%. TSLO reports NAN during problem periods. Resetting sensor during issue did not resolve problem. Please replace sensor.
HECT Wind Speed at 2 m 2020-12-01 20:30:41 Mice nest found inside enclosure. Technician confirmed damage to WS2M cable. Damage patched with electrical tape. No problems found in data at this time. Replace sensor. Will increase priority if data become affected.
EVAX 229L Delta Temperature at 60 cm 2020-09-21 21:27:53 Both soil temperature and soil moisture frequently report -7999. Please replace sensor.